OpenStudy (anonymous):

Find an equation for the tangent to the curve at the given point: f(x)=4(sqrt(x))-x+3, (16,3)

OpenStudy (asnaseer):

do you know how to differentiate?

OpenStudy (anonymous):

\[f(x)=4\sqrt{x}-x+3\]

OpenStudy (anonymous):

Yes but not so well with radicals

OpenStudy (asnaseer):

remember that \(\sqrt{x}\) can be written as \(\displaystyle x^{\frac{1}{2}}\)

OpenStudy (asnaseer):

does that help?

OpenStudy (anonymous):

so 2/sqrt(x)

OpenStudy (asnaseer):

that is the first term - yes

OpenStudy (anonymous):

\[d/dx=2/\sqrt{x}-1\]

OpenStudy (asnaseer):

yup, although usually written out as:\[f'(x)=\frac{2}{\sqrt{x}}-1\]

OpenStudy (anonymous):

so what do I do with coordinates (16,3)? Just substitute?

OpenStudy (asnaseer):

the differential will give you the value of the slope of the tangent at any point

OpenStudy (asnaseer):

so, since the differential involves terms only containing x, you would substitute x=16 to get the value of the slope of the tangent line

OpenStudy (anonymous):

so 1/2 is the slope and 3 is the y intercept so the answer should be y=1/2x+3

OpenStudy (asnaseer):

that doesn't seem right to me

OpenStudy (anonymous):

oohhhso 2/4x-16+3 the answer is 1/2x-11??

OpenStudy (asnaseer):

lets do this one step at a time to avoid mistakes. first what do you get for the value of the slope at (16,3) ?

OpenStudy (anonymous):

2/sqrt(16)x =2/4x =1/2x

OpenStudy (anonymous):

Or (1/2)x

OpenStudy (asnaseer):

no - remember the equation that tells us what the slope of the tangent line will be at any point is given by:\[f'(x)=\frac{2}{\sqrt{x}}-1\]so substitute x=16 into this equation and what do you get for the value of f'(x)?

OpenStudy (anonymous):

-1/2

OpenStudy (asnaseer):

correct, so now you know the equation of the tangent line is:\[y=-\frac{1}{2}x +c\]

OpenStudy (asnaseer):

next, we know that this line must also pass through the point (16,3). So use this to calculate the value for c.

OpenStudy (anonymous):

so I need to solve -(1/2)(16)+c=0?

OpenStudy (anonymous):

or =3?

OpenStudy (asnaseer):

not =0, but =3

OpenStudy (anonymous):

ok give me one sec to get this

OpenStudy (asnaseer):

remember it passes through the point (16,3) => x=16, y=3

OpenStudy (anonymous):

that cant be right...-(1/2)(16)=-8 c=3-8 =-5 That is not an option. It has to be 3, -11, or 11

OpenStudy (anonymous):

nevermind 11 :)

OpenStudy (asnaseer):

you made a mistake. remember the equation is:\[y=-\frac{1}{2}x +c\]

OpenStudy (anonymous):

I think you know what I did :D

OpenStudy (asnaseer):

:)

OpenStudy (anonymous):

Thank you so much for you help

OpenStudy (asnaseer):

yw :)